# Radiative corrections and Monte Carlo tools for low-energy hadronic cross sections in $e^{+}e^{-}$ collisions

### v2

@@ 2 Experiments at electron–positron colliders
Many experiments have contributed to the measurements of the $e^{+}e^{-}\to\pi^{+}\pi^{-}$ channel over the years, as shown in Figure 1. Information about two-pion cross section measurements is collected in the database
@@ 3.1 Fixed-order QED
over the two-body phase space. Here we introduced the notation $\mathcal{M}_{xx}^{(n)}$ to include all terms of the squared matrix elements with a power $\mathcal{O}(\alpha^{n})$ relative to the LO term. After removing the UV singularities in (3.3) through renormalisation (usually in the on-shell scheme) there are still IR singularities. They cancel for physical (IR-safe) observables if the virtual corrections are combined with real corrections. The latter are obtained by integrating
@@
Keeping finite fermion masses, there are no collinear singularities in QED. Instead, collinear logarithms $L_{c}$ appear. However, there are still soft singularities. They can be regularised either by introducing a fictitious photon mass $m_{\gamma}$ or by using dimensional regularisation also for the IR case. The regularised singularities then appear as $\log(Q^{2}/m_{\gamma}^{2})$ or $1/\epsilon$ poles in intermediate expressions. The extraction of the IR singularities of the real corrections can be done in a process- and observable-independent way, using either the slicing method or a subtraction method. A short description of the two can be found, e.g., in [112]. Once real and virtual corrections are combined, it is safe to remove the regulator $m_{\gamma}\to 0$ or $\epsilon\to 0$. While not strictly necessary, typically the slicing method is combined with photon-mass regularisation, whereas the subtraction method is used with dimensional regularisation.
@@
and integrate (3.5) and (3.6) over the two-body and three-body phase space, respectively, to obtain $\mathrm{d}\sigma^{(1)}_{mm}(q_{e}^{4}\,q_{m}^{2})$. Analogously, we can obtain $\mathrm{d}\sigma^{(1)}_{mm}(q_{e}^{2}\,q_{m}^{4})$, the final-state NLO corrections to $e^{+}\,e^{-}\to\mu^{+}\,\mu^{-}$. Finally, there are the mixed corrections consisting of $\mathrm{d}\sigma^{(1)}_{mm}(q_{e}^{3}\,q_{m}^{3})$ (interference between emission of initial state and final state) and $\mathrm{d}\sigma_{mm}^{(1)}(q_{e}^{2}\,q_{m}^{2}\,\Pi^{(1)})$ (VP corretions). The sum of these four parts is denoted by
@@ 3.4.1 Strategies for radiative corrections to low-energy hadronic amplitudes
Refer to caption $\times$
Figure 7: LO contribution to the $e^{+}e^{-}\to\pi^{+}\pi^{-}$ cross section.
@@ 4 Generators and integrators
The seven codes discussed during Phase I are AfkQed, Babayaga@NLO, KKMCee, MCGPJ, McMule, Phokhara, and Sherpa. A list of responsible contact persons for each code can be found on the website at
@@
Other codes may be added in Phase II.
@@ 4.1 AfkQed
HVP None, leptonic only [278], Customisable
leptonic [278] + hadronic, form factor
@@ 4.3 KKMC
Effects due to photon emission from incoming beams and outgoing fermions are calculated in QED up to second order in $\alpha_{\text{QED}}$, including all interference effects, within Coherent Exclusive Exponentiation (CEEX) [155, 290], which is based on Yennie–Frautschi–Suura exponentiation [153], see Section 3.2.3. CEEX not only treats infrared cancellations and the description of soft photons correctly to infinite order, but also includes the correct description of QED interferences and narrow resonances. Contrary to the standard YFS approach, in which higher-order corrections are effected at the level of the amplitude squared, the CEEX scheme is devised in terms of spin amplitudes and can therefore account for ISR-FSR interferences. A detailed description of CEEX and the older EEX scheme can be found in [155, 291]. In addition to higher order QED corrections the weak corrections are provided by Dizet library [292], and polarised $\tau$ decays are included using the TAUOLA programme [293].
@@ 4.5 McMule
$\mathcal{M}_{n+\gamma}^{(1)}=\bigl{(}{\cal E}+{\cal D}+{\cal S}^{(1)}\bigr{)}\,\mathcal{M}_{n}^{(1)}+\mathcal{O}(E_{\gamma}^{0})\,,$ (4.5)
@@ 4.6 Phokhara
The treatment of hadrons in the final state, in particular pions, is carried out under the approximation of the pions being point-like particles described by sQED (see Figure 2). This approach, not being rigorous, is augmented by the use of a non-perturbative form factor. For further details, we refer the reader to [323, Figure 2.10].
The decomposition between perturbative and non-perturbative contributions to the pion production cross section (1.2a) motivated the addition of initial-stae radiative corrections as a first attempt in Phokhara to simulate emission of photons at large-angle ($\theta\gg m_{e}^{2}/s$[324, 325, 107], and small-angle ($\theta\ll m_{e}^{2}/s$[326] regions. To further improve theoretical predictions, for final-state radiation off the pions, Phokhara adopted the F×$\times$×sQED approach (see Section 3.4.1). The accuracy of this approach is under scrutiny. Using these methods, Phokhara includes the emission of one photon from the initial state and one photon from the final state, requiring one of the photons to be hard [258, 248]. The relevant Feynman diagrams can be seen in the right column of Figure 13 and Figure 15.
@@ 4.7 Sherpa
A key advantage of the EEX method is its relatively straightforward implementation, as well as the ease with which one can verify the behaviour of these expressions in specific limits, such as soft or collinear limits. Additionally, constructing terms independently for ISR and FSR contributions simplifies the automation and implementation of ISR effects, particularly for lepton colliders (such as $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$). However, due to the potential complexity of the final states, automating FSR contributions is challenging and generally requires handling each case individually. The EEX corrections we have included were computed in [340, 341, 130, 342, 343, 143, 344], and explicit expressions for the infrared-subtracted terms can be found in [290]. We summarise these corrections in Table 9.
@@ 5 Monte Carlo comparisons
In this section we present results obtained with the Monte Carlo codes described in Section 4. To this end, we define a set of scenarios, i.e., simplified setups consisting of acceptance cuts, reminiscent of the experiments described in Section 2. The purpose is not to precisely match the experimental analyses carried out. Rather, we want to provide simplified but still realistic phenomenological scenarios that can be used to validate the codes and assess the importance of the various contributions that contribute to differential cross sections. The source codes that have been used to obtain these results can be found at
@@ 5.1.1 Vacuum polarisation
All packages mentioned above include at least the NLO term in the leptonic part of the VP. This is more than sufficient compared to the precision level of the hadronic part. The NSK tabulation of the leptonic $\Pi_{\ell}(s)$ was calculated numerically through the integral from the total $e^{+}e^{-}\to\ell^{+}\ell^{-}\gamma$ cross section. This includes the LO FSR correction [249] and the additional enhancement on the threshold from the Coulomb final-state interaction. (the $\Pi_{\ell}(s)$ by the dispersive integral from the such cross section without the Sommerfeld–Gamow–Sakharov factor gives numerically identical result to the one-two-loop analytical formula of the VP in the paper [109]).
@@
In order to disentangle effects from different versions of VP from other differences, we use [169] as default for the hadronic VP for the results presented in this section. However, the procedure of resummation of VPC – or the lack thereof – differs between various codes. These differences are typically beyond NLO, but can be numerically significant, in particular near resonances.
@@ Acknowledgements
The work in this report is supported by the Leverhulme Grant LIP-2021-014, the STFC (grants ST/P001246/1, ST/T000988/1, and ST/X000699/1), the DFG through the funds provided to the Sino-German Collaborative Research Center TRR110 “Symmetries and the Emergence of Structure in QCD” (DFG Project-ID 196253076 – TRR 110), the SNSF (Project Nos. 200020_200553, 200020_207386, PCEFP2_181117, PCEFP2_194272, and TMSGI2_211209), the Italian Ministero dell’Università e Ricerca (MUR) and European Union - Next Generation EU through the research grants 2022BCXSW9, 20225X52RA, and 2022ENJMRS under the programme PRIN 2022, the European Union STRONG2020 project under Grant Agreement Number 824093, the Polish National Science Centre through the Grants Nos. 2019/35/O/ST2/02907 and 2023/50/A/ST2/0022, and by the Priority Research Area Digiworld under the programme “Excellence Initiative – Research University” at the Jagiellonian University in Krakow.

### v3

@@ 2 Experiments at electron–positron colliders
Many experiments have contributed to the measurements of the $e^{+}e^{-}\to\pi^{+}\pi^{-}$ channel over the years, as shown in Figure 1. Their use for a data-driven determination of the HVP is described in detail in [6]. Information about two-pion cross-section measurements is collected in the database
@@ 3.1 Fixed-order QED
over the two-body phase space. Here we introduced the notation $\mathcal{M}_{xx}^{(n)}$ to include all terms of the squared matrix elements with a power $\mathcal{O}(\alpha^{n})$ relative to the LO term. After removing the UV singularities in (3.3) through renormalisation (usually in the on-shell scheme) there are still IR singularities. As discussed below, they cancel for physical (IR-safe) observables if the virtual corrections are combined with real corrections. The latter are obtained by integrating
@@
Keeping finite fermion masses, there are no collinear singularities in QED. Instead, the already mentioned collinear logarithms of the form $L_{c}=\log(Q^{2}/m_{\ell}^{2})$ arise in the angular integration of radiation degrees of freedom around the emitting fermion lines. For large gaps between the two scales $Q^{2}$ and $m_{\ell}^{2}$, as is typically the case for $\ell=e$, such logarithms enhance the higher-order terms of the perturbative series at fixed order. In Sections 3.2.1 and 3.2.2 the approximations able to resum the collinear logarithms to all orders are discussed.
Even with finite fermion masses there are still soft singularities. They can be regularised either by introducing a fictitious photon mass $m_{\gamma}$ or by using dimensional regularisation also for the IR case. The regularised singularities then appear as $\log(Q^{2}/m_{\gamma}^{2})$ or $1/\epsilon$ poles in intermediate expressions. The extraction of the IR singularities of the real corrections can be done in a process- and observable-independent way, using either the slicing method or a subtraction method. A short description of the two can be found, e.g., in [112]. Once real and virtual corrections are combined, it is safe to remove the regulator $m_{\gamma}\to 0$ or $\epsilon\to 0$. While not strictly necessary, typically the slicing method is combined with photon-mass regularisation, whereas the subtraction method is used with dimensional regularisation.
The core property that lies at the heart of the IR cancellation is the fact that it is impossible in principle to distinguish an electron (or any charged lepton) from an electron accompanied by a cloud of arbitrarily soft photons. Even the best detector has a finite resolution below which photons remain undetected. From a more formal point of view, the IR divergences of the scattering matrix stem from the violation of a basic assumption of scattering theory: due to the long-range nature of the Coulomb interaction, the electrons are never free, not even in the asymptotic past or future.
From a practical point of view, this is acceptable as long as the observables considered are defined accordingly. They must not depend on whether or not additional arbitrarily soft photons are emitted. For such so-called IR-safe observables the cancellation between real and virtual IR singularities is guaranteed to all orders in perturbation theory. However, it is often the case that as a remnant of this cancellation there are (large) soft logarithms. Indeed, if the observable is defined such that the real emission of additional photons is restricted to have energies $\Delta{E}$, there are logarithms of the form $\alpha^{n}\log^{n}(Q^{2}/\Delta{E}^{2})$. The precise form of the soft logarithms depends on the definition of the observable. For small $\Delta{E}$, the soft logarithms can be numerically enhanced, potentially leading to a bad convergence of the perturbative expansion. The inclusion of these logarithms beyond fixed order is discussed in Section 3.2.3.
@@
and integrate (3.5) and (3.6) over the two-body and three-body phase space, respectively, to obtain $\mathrm{d}\sigma^{(1)}_{mm}(q_{e}^{4}\,q_{m}^{2})$. Analogously, we can obtain $\mathrm{d}\sigma^{(1)}_{mm}(q_{e}^{2}\,q_{m}^{4})$, the final-state NLO corrections to $e^{+}\,e^{-}\to\mu^{+}\,\mu^{-}$. Finally, there are the mixed corrections consisting of $\mathrm{d}\sigma^{(1)}_{mm}(q_{e}^{3}\,q_{m}^{3})$ (interference between emission of initial state and final state) and $\mathrm{d}\sigma_{mm}^{(1)}(q_{e}^{2}\,q_{m}^{2}\,\Pi^{(1)})$ (VP corrections). The sum of these four parts is denoted by
@@ 3.4.2 Structure-dependent radiative corrections in $\boldsymbol{e^{+}e^{-}\to\pi^{+}\pi^{-}}$
Refer to caption $\times$
Figure 7: LO contribution to the $e^{+}e^{-}\to\pi^{+}\pi^{-}$ cross section.
@@ 4 Generators and integrators
The seven codes discussed during Phase I are AfkQed, Babayaga@NLO, KKMCee, MCGPJ, McMule, Phokhara, and Sherpa. These codes have been developed by groups and authors that are not necessarily part of this community effort. While the responsibility of the codes clearly lies with the original authors, we provide list of responsible contact persons for each code on the website at
@@
This list can serve as a first point of contact for potential users of the codes. Other codes may be added in Phase II.
@@ 4.1 AfkQed
HVP None, leptonic only [278], F×$\times$×sQED
leptonic [278] + hadronic, Customisable VFF
@@ 4.3 KKMC
Effects due to photon emission from incoming beams and outgoing fermions are calculated in QED up to second order in $\alpha_{\text{QED}}$, including all interference effects, within Coherent Exclusive Exponentiation (CEEX) [155], which is based on Yennie–Frautschi–Suura exponentiation [153], see Section 3.2.3. CEEX not only treats infrared cancellations and the description of soft photons correctly to infinite order, but also includes the correct description of QED interferences and narrow resonances. Contrary to the standard YFS approach, in which higher-order corrections are effected at the level of the amplitude squared, the CEEX scheme is devised in terms of spin amplitudes and can therefore account for ISR-FSR interferences. A detailed description of CEEX and the older EEX scheme can be found in [155, 290]. In addition to higher order QED corrections the weak corrections are provided by Dizet library [291], and polarised $\tau$ decays are included using the TAUOLA programme [292].
@@ 4.5 McMule
$\mathcal{M}_{n+\gamma}^{(1)}=\bigl{(}{\cal E}+{\cal D}\bigr{)}\,\mathcal{M}_{n}^{(1)}+{\cal S}^{(1)}\,\mathcal{M}_{n}^{(0)}+\mathcal{O}(E_{\gamma}^{0})\,,$ (4.5)
@@ 4.6 Phokhara
The treatment of hadrons in the final state, in particular pions, is carried out under the approximation of the pions being point-like particles described by sQED (see Figure 2). This approach, not being rigorous, is augmented by the use of a non-perturbative form factor. For further details, we refer the reader to [322, Figure 2.10].
The decomposition between perturbative and non-perturbative contributions to the pion production cross section (1.2a) motivated the addition of initial-state radiative corrections as a first attempt in Phokhara to simulate emission of photons at large-angle ($\theta\gg m_{e}^{2}/s$[323, 324, 107], and small-angle ($\theta\ll m_{e}^{2}/s$[325] regions. To further improve theoretical predictions, for final-state radiation off the pions, Phokhara adopted the F×$\times$×sQED approach (see Section 3.4.1). The accuracy of this approach is under scrutiny. Using these methods, Phokhara includes the emission of one photon from the initial state and one photon from the final state, requiring one of the photons to be hard [258, 248]. The relevant Feynman diagrams can be seen in the right column of Figure 13 and Figure 15.
@@ 4.7 Sherpa
A key advantage of the EEX method is its relatively straightforward implementation, as well as the ease with which one can verify the behaviour of these expressions in specific limits, such as soft or collinear limits. Additionally, constructing terms independently for ISR and FSR contributions simplifies the automation and implementation of ISR effects, particularly for lepton colliders (such as $e^{+}e^{-}$ or $\mu^{+}\mu^{-}$). However, due to the potential complexity of the final states, automating FSR contributions is challenging and generally requires handling each case individually. The EEX corrections we have included were computed in [339, 340, 130, 341, 342, 143, 343, 344, 345], and explicit expressions for the infrared-subtracted terms can be found in [155]. We summarise these corrections in Table 9.
@@ 5 Monte Carlo comparisons
In this section we present results obtained with the Monte Carlo codes described in Section 4. To this end, we define a set of scenarios, i.e., simplified setups consisting of acceptance cuts, reminiscent of the experiments described in Section 2. The purpose is not to precisely match the experimental analyses carried out. Rather, we want to provide simplified but still realistic phenomenological scenarios that can be used to validate the codes and assess the importance of the various contributions to differential cross sections. As we will see, the impact and importance of these contributions is strongly dependent on the observables. Thus, generic statements are difficult to make and should be treated with care.
The source codes that have been used to obtain these results can be found at
@@ 5.1.1 Vacuum polarisation
All packages mentioned above include at least the NLO term in the leptonic part of the VP. This is more than sufficient compared to the precision level of the hadronic part. The NSK tabulation of the leptonic $\Pi_{\ell}(s)$ was calculated numerically through the integral from the total $e^{+}e^{-}\to\ell^{+}\ell^{-}\gamma$ cross section. This includes the LO FSR correction [249] and the additional enhancement on the threshold from the Coulomb final-state interaction. (the $\Pi_{\ell}(s)$ by the dispersive integral from such cross sections without the Sommerfeld–Gamow–Sakharov factor gives numerically identical results to the one-two-loop analytical formula of the VP in the paper [109]). Similar methods have been used to describe contributions of low-virtuality photons to jets at LHC [348].
@@
In order to disentangle effects from different versions of VP from other differences, we use [169] as default for the hadronic VP for the results presented in this section. However, the procedure of resummation of VPC – or the lack thereof – differs between various codes. These differences are typically beyond NLO, but can be numerically significant, in particular near resonances. The precise treatment of the VPC in the codes is described in Section 4 and in the following comparisons we will highlight their impact for the observables under consideration.
@@ Acknowledgements
The work in this report is supported by the Leverhulme Trust LIP-2021-014, the STFC (grants ST/P001246/1, ST/T000988/1, and ST/X000699/1), the DFG through the funds provided to the Sino-German Collaborative Research Center TRR110 “Symmetries and the Emergence of Structure in QCD” (DFG Project-ID 196253076 – TRR 110), the SNSF (Project Nos. 200020_200553, 200020_207386, PCEFP2_181117, PCEFP2_194272, and TMSGI2_211209), the Italian Ministero dell’Università e Ricerca (MUR) and European Union - Next Generation EU through the research grants 2022BCXSW9, 20225X52RA, and 2022ENJMRS under the programme PRIN 2022, the European Union STRONG2020 project under Grant Agreement Number 824093, the Polish National Science Centre through the Grants Nos. 2019/35/O/ST2/02907 and 2023/50/A/ST2/0022, and by the Priority Research Area Digiworld under the programme “Excellence Initiative – Research University” at the Jagiellonian University in Krakow. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no. PLG/2024/017001.